OpenStudy (anonymous):

What volume of a .200 M solution of K2SO4 solution contains 85.6 g of K2SO4?

OpenStudy (frostbite):

Hey @johnybgood199 and welcome to \(\LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study!}}\) In order to solve this problem we need primarily two equations:\[\Large C=\frac{ n }{ V }\]and \[\Large n=\frac{ m }{ M }\] Where: \(C\) is the concentration [M] \(V\) is the volume [L] \(n\) is the amount of substance [mol] \(m\) is the mass [g] \(M\) is the molar mass [g/mol] You know how to obtain the molar mass for\(\ \sf K_{2}SO_{4}\)?

OpenStudy (anonymous):

Yes 174.259 g/mol

OpenStudy (anonymous):

2.456L

OpenStudy (frostbite):

Perfect, then you just put into the equations and solve for the volume \(V\)

OpenStudy (anonymous):

85.6g * 1mol / 174.259g = 0.491 mol V = 0.491mol/0.2mol

OpenStudy (anonymous):

V = 2.456 L

OpenStudy (anonymous):

is that right?

OpenStudy (frostbite):

It sure is, but in your last calculations change the mol to M. Also I suggest you give the final answer with only 3 significant figures as there are only given 3 in the question. Else well done!

OpenStudy (frostbite):

So it become mol/M

OpenStudy (frostbite):

And ones again welcome to OpenStudy!

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

Can I ask you another question?